OpenStudy (anonymous):

taylor series approximation of sin x ,x=pi/3

OpenStudy (anonymous):

\[Given f(x)=\sin x ,taylor series approximations if x= \pi \div3,order 0 \to 7.\]

OpenStudy (anonymous):

f(x)=f(a)+f'(a)(xa)+[f"(a)/2! ](x-a)^2 +[f"'(a)/3!](x-a)^3 +[f^(n)/n!](x-a)^n+..... [sqrt3/2]+[(x-pi/3)/2]-[sqrt3(x-pi/3)^2]/4 -[(x-pi/3)^3]/12 +[(x-pi/3)^4]/16sqrt3 +[(x-pi/3)^5]/240 + 0[(x-pi/3)^6] answer